OpenStudy (anonymous):

1.If I know the lengths of two sides of a right triangle, how do I find the third?

OpenStudy (anonymous):

is it right angled

OpenStudy (anonymous):

you have to know an angle to find the third side

OpenStudy (anonymous):

is there a step by step process to finding this?

OpenStudy (anonymous):

if it is right triangle then \[x^2+y^2=r^2\] 

OpenStudy (anonymous):

so if i have a triangle ABC. A=6 B=9 and we must find the length of C. i must multiply AxA and add that to the product BxB to find C squared. and to find C i must use the radical sign to undo C squared basically. correct me if im wrong? i think this is the pythagorean theorum

Parth (parthkohli):

This is known as the Pythagorean Theorem.

Parth (parthkohli):

\[a^2 + b^2 = c^2\]If a,b are two legs and c is the hypotenuse(longest side), then the above is true in a right triangle.

OpenStudy (anonymous):

okay thank you!(:

Parth (parthkohli):

I like examples, so I'd give one: Find \(x\). We know that 5 is the longest side because it is opposite 90. Also, \(x\) and 3 are legs.